OpenStudy (oswaldmurphy):

The lengths of two sides of a triangle are shown below: Side 1: 3x2 − 2x − 1 Side 2: 9x + 2x2 − 3 The perimeter of the triangle is 5x3 + 4x2 − x − 3. Part A: What is the total length of the two sides, 1 and 2, of the triangle? (4 points) Part B: What is the length of the third side of the triangle? (4 points) Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer. (2 points)

OpenStudy (oswaldmurphy):

This is a test! no direct answers please! just help me understand them. Part A you add side 1 and 2 together. But when you add 3x^2 and 2x^2 what is that? Please help! @TheSmartOne

OpenStudy (oswaldmurphy):

Im not asking for an answer.. Just help me understand it. I would do this to a teacher but its a online course. I am blazing through this so please.. I have done almost 2 semesters over 2 months and Im going in 8th grade and this is 9th, 10th grade work.

OpenStudy (sainath):

pls help me at mine

OpenStudy (oswaldmurphy):

You cant go to other questions and ask for help...

TheSmartOne (thesmartone):

basically 3x^2 + 2x^2 = x^2(3 + 2) = x^2(5) = 5x^2 you can add them because they are like terms, they both have x^2 and you just add the coefficients you have 3 cookies and you get 2 more. How many do you have? 5 cookies just like that 3x^2 + 2x^2 = 5x^2

OpenStudy (oswaldmurphy):

@TheSmartOne Ok. I got A done. I need help with B. Do I subtract the perimeter from the Answer for A?

TheSmartOne (thesmartone):

Correct :)

TheSmartOne (thesmartone):

well actually you do the opposite of that you subtract the answer you got for A from the perimeter but I'm sure that's what you meant xD

OpenStudy (oswaldmurphy):

Is this the correct answer? 11x + 4x^2 - 7

TheSmartOne (thesmartone):

Nope